# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ Appendix E Auxiliaries for the ergodicity of i-SIR
It holds for all $\lambda,\lambda^{\prime}\in[1,\infty)$ and $x\in\mathbb{S}$, that
$d_{\mathrm{tv}}\big(P_{\lambda}(x,\;\cdot\;),P_{\lambda^{\prime}}(x,\;\cdot\;)\big)\leq 2|\lambda-\lambda^{\prime}|.$
@@
Let $N,N^{\prime}\in\mathbb{N}_{+}$. Let $N\leq\lambda<\lambda^{\prime}\leq N+1$, then
$d_{\mathrm{tv}}(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;))$ $=\sup_{A}|P_{\lambda^{\prime}}(x,A)-P_{\lambda}(x,A)|$
$=\sup_{A}|(\lambda^{\prime}-\lambda)P_{N+1}(x,A)-(\lambda-\lambda^{\prime})P_{N}(x,A)|$
$=|\lambda-\lambda^{\prime}|\sup_{A}|P_{N}(x,A)+P_{N+1}(x,A)|\leq 2|\lambda^{\prime}-\lambda|.$
We get by triangle inequality that if $N<N^{\prime}$, then
$d_{\mathrm{tv}}(P_{N^{\prime}}(x,\;\cdot\;),P_{N}(x,\;\cdot\;))\leq\sum_{i=N}^{N^{\prime}-1}d_{\mathrm{tv}}(P_{i+1}(x,\;\cdot\;),P_{i}(x,\;\cdot\;))=2(N^{\prime}-N)$
and if $N<\lambda\leq N+1\leq N^{\prime}<\lambda^{\prime}\leq N^{\prime}+1$, then
$d_{\mathrm{tv}}(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;))$
$\leq d_{\mathrm{tv}}(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{N^{\prime}}(x,\;\cdot\;))+d_{\mathrm{tv}}(P_{N^{\prime}}(x,\;\cdot\;),P_{N+1}(x,\;\cdot\;))+d_{\mathrm{tv}}(P_{N+1}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;))$
$\leq 2(\lambda^{\prime}-N^{\prime})+2(N^{\prime}-N-1)+2(N+1-\lambda)=2(\lambda^{\prime}-\lambda).$
Thus if $\lambda,\lambda^{\prime}\in[1,\infty)$, then
$d_{\mathrm{tv}}(P_{\lambda}^{\prime}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;))\leq 2|\lambda^{\prime}-\lambda|.\qed$
@@
which also holds for $N=1$, since $P_{1}(z_{1},A)\geq 0$. If $\lambda\geq 1$, $N\leq\lambda<N+1$, where $N\in\mathbb{N}_{+}$, then
@@
$P_{N}(z_{1},A)$ $=\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\in A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$=1-\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$=1-(N-1)\int_{\mathbb{X}^{N-1}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{w(z_{1})+w(z_{2})+N-2}q(\mathrm{d}z_{2})$
$\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{2\hat{w}+N-2}q(\mathrm{d}z_{2})=1-(N-1)\frac{\pi(A^{c})}{2\hat{w}+N-2}$
$=\pi(A)+\pi(A^{c})\frac{2\hat{w}-1}{2\hat{w}+N-2},$
where the first inequality follows from Jensen’s inequality. The case $N=1$ is clear. By Lemma 45 in Appendix D it follows, that for all $\lambda\geq 1$
@@
By the first lower bound in Lemma 48 we have for all $x\in\mathbb{S}$ and $\lambda\geq s$ a constant $c=\frac{s-1}{2\hat{w}+s-1}\in(0,1)$ such that
@@ Appendix F Cost function
$u(\lambda)-c(\lambda)=\frac{d(a+b\lambda)}{\lambda-1}\xrightarrow{\lambda\to\infty}bd$
@@ Appendix G Finite discrete case for i-SIR
Here we will use the same notations and assumptions as in Section 6. Estimating the asymptotic variance of i-SIR can be challenging. In the finite discrete case asymptotic variance can be solved via the formula in Lemma 39 in Appendix A, since the transition probability matrix $P_{\lambda}$ as well as the eigenvalues and eigenvectors can be solved, even tough this can be extremely slow for large $\lambda$, when $\mathbb{S}$ has many states. However in the finite case transition probability matrix $P_{N}$ can be expressed with the help of multinomial distribution:
@@
Since $(X_{i})$ is Markov chain we obtain the case for $U=\emptyset$. The case $U\neq\emptyset$ follows similarly, when we have instead $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}),q(U))$, $v\in\mathbb{R}^{n+1}$,
@@
If $(\mathbb{S})=2$, i.e. $\mathbb{S}=\{s_{1},s_{2}\}$, it holds for $f\in L_{0,+}^{2}(\pi)$
@@
We have $P_{\lambda}$ is $\pi$-reversible by Lemma 26. We can solve eigenvalues $\lambda_{1}=1$, $\lambda_{2}=P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})-1$ and eigenvectors $v_{1}=(1,1)$, $v_{2}=(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}},\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}})$ to match conditions in Lemma 39 in Appendix A and thus
$\mathrm{var}(P_{\lambda},f)$ $=\langle v_{2},f\rangle_{\pi}^{2}\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-(P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\}))}$
@@
$\langle v_{2},f\rangle_{\pi}^{2}$ $=\bigg(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}}f(s_{1})\pi(\{s_{1}\})+\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}}f(s_{2})\pi(\{s_{2}\})\bigg)^{2}$
@@
$\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{1}\})+\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{2}\})-1=-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})$
$\quad=-\big(\pi(\{s_{1}\})+\pi(\{s_{2}\})\big)\big(\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})+\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})\big)$
@@
$H_{f}(\lambda)$ $=\frac{1+\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{1}\})+\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{2}\})-2\pi(\{s_{1}\})^{2}-2\pi(\{s_{2}\})^{2}]}{1-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{1}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{2}\})}\mathrm{var}_{\pi}(f)$
$=\frac{2-[P_{\lambda}(s_{1},\{s_{2}\})+P_{\lambda}(s_{2},\{s_{1}\})]}{P_{\lambda}(s_{1},\{s_{2}\})+P_{\lambda}(s_{2},\{s_{1}\})}\mathrm{var}_{\pi}(f)=\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-[P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})]}\mathrm{var}_{\pi}(f)$

### v2

@@ Appendix E Auxiliaries for the ergodicity of i-SIR
Let $\in\mathbb{N}_{+}, M<NM<N$, then
$|P_{M}(x,A)-P_{N}(x,A)|\leq\frac{N-M}{N-1}.$
@@
Let us denote $z_{1}=x$, $B_{n}=\sum_{i=1}^{n}w(z_{i})\mathbf{1}\left(z_{i}\in A\right)$ and $C_{n}=\sum_{i=1}^{n}w(z_{i})$, then
$|P_{M}(x,A)-P_{N}(x,A)|$
$\quad=\bigg|\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{B_{M}C_{N}-B_{N}C_{M}}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\bigg|$
$\quad=\bigg|\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{i=1}^{M}\sum_{j=M+1}^{N}w(z_{i})w(z_{j})[\mathbf{1}\left(z_{i}\in A\right)-\mathbf{1}\left(z_{j}\in A\right)]}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\bigg|$
$\quad\leq\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{i=1}^{M}\sum_{j=M+1}^{N}w(z_{i})w(z_{j})}{C_{M}C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$\quad=\int_{\mathbb{W}_{M}\times\mathbb{X}^{N-M}}\frac{\sum_{j=M+1}^{N}w(z_{j})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)=\int_{\mathbb{W}_{N}}\frac{\sum_{j=M+1}^{N}w(z_{j})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)$
$\quad=\frac{N-M}{N-1}\int_{\mathbb{W}_{N}}\frac{\sum_{i=2}^{N}w(z_{i})}{C_{N}}\bigg(\prod_{n=2}^{N}q(\mathrm{d}z_{n})\bigg)\leq\frac{N-M}{N-1},$
where the last equality follows since $z_{2},\dots,z_{N}$ are exchangeable under the integral. ∎
Lemma 48.
It holds for all $\lambda,\lambda^{\prime}\in[1,\infty)$ and $x\in\mathbb{S}$ that
$d_{\mathrm{tv}}\big(P_{\lambda}(x,\;\cdot\;),P_{\lambda^{\prime}}(x,\;\cdot\;)\big)\leq\frac{6|\lambda-\lambda^{\prime}|}{\max\{\lambda,\lambda^{\prime}\}}.$
Proof.
Let $N,N^{\prime}\in\mathbb{N}_{+}$. If $N\leq\lambda<\lambda^{\prime}\leq N+1$, then $1/N\leq 2/\lambda^{\prime}$ and
$d_{\mathrm{tv}}\big(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;)\big)$
$\quad=\sup_{A}|P_{\lambda^{\prime}}(x,A)-P_{\lambda}(x,A)|=\sup_{A}|(\lambda-\lambda^{\prime})P_{N}(x,A)+(\lambda^{\prime}-\lambda)P_{N+1}(x,A)|$
$\quad=|\lambda^{\prime}-\lambda|\sup_{A}|P_{N+1}(x,A)-P_{N}(x,A)|\leq\frac{1}{N}|\lambda^{\prime}-\lambda|,$
where the last inequality follows from Lemma 47. If $N<\lambda\leq N+1\leq N^{\prime}<\lambda^{\prime}\leq N^{\prime}+1$, then by the previous inequality, Lemma 47 and triangle inequality
$d_{\mathrm{tv}}\big(P_{N^{\prime}}(x,\;\cdot\;),P_{N}(x,\;\cdot\;)\big)$
$\quad\leq d_{\mathrm{tv}}\big(P_{\lambda^{\prime}}(x,\;\cdot\;),P_{N^{\prime}}(x,\;\cdot\;)\big)+d_{\mathrm{tv}}\big(P_{N^{\prime}}(x,\;\cdot\;),P_{N+1}(x,\;\cdot\;)\big)+d_{\mathrm{tv}}\big(P_{N+1}(x,\;\cdot\;),P_{\lambda}(x,\;\cdot\;)\big)$
$\quad\leq|\lambda^{\prime}-N^{\prime}|\frac{1}{N^{\prime}}+|N^{\prime}-(N+1)|\frac{1}{N^{\prime}-1}+|N+1-\lambda|\frac{1}{N}\leq\frac{\lambda^{\prime}-(N+1)}{N^{\prime}-1}+\frac{N+1-\lambda}{N}$
$\quad=\frac{N(\lambda^{\prime}-\lambda)+(N^{\prime}-1-N)(N+1-\lambda)}{N(N^{\prime}-1)}\quad\leq\frac{N(\lambda^{\prime}-\lambda)+N^{\prime}-(N+1)}{N(N^{\prime}-1)}$
$\quad\leq\frac{N(\lambda^{\prime}-\lambda)+\lambda^{\prime}-\lambda}{N(N^{\prime}-1)}=\frac{(N+1)(\lambda^{\prime}-\lambda)}{N(N^{\prime}-1)}\leq\frac{2(\lambda^{\prime}-\lambda)}{N^{\prime}-1}\leq\frac{6(\lambda^{\prime}-\lambda)}{N^{\prime}+1}\leq\frac{6(\lambda^{\prime}-\lambda)}{\lambda^{\prime}}.$
Therefore, we obtain the claim. ∎
@@
which also holds for $N=1$, since $P_{1}(z_{1},A)\geq 0$. If $\lambda\geq 1$, $N\in\mathbb{N}_{+}$, $N\leq\lambda<N+1$, then
@@
$P_{N}(z_{1},A)$
$\quad=\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\in A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})=1-\int_{\mathbb{X}^{N-1}}\frac{\sum_{i=1}^{N}w(z_{i})\mathbf{1}\left(z_{i}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$\quad=1-(N-1)\int_{\mathbb{X}^{N-1}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{\sum_{k=1}^{N}w(z_{k})}\prod_{n=2}^{N}q(\mathrm{d}z_{n})$
$\quad\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{w(z_{1})+w(z_{2})+N-2}q(\mathrm{d}z_{2})\leq 1-(N-1)\int_{\mathbb{X}}\frac{w(z_{2})\mathbf{1}\left(z_{2}\notin A\right)}{2\hat{w}+N-2}q(\mathrm{d}z_{2})$
$\quad=1-(N-1)\frac{\pi(A^{c})}{2\hat{w}+N-2}=\pi(A)+\pi(A^{c})\frac{2\hat{w}-1}{2\hat{w}+N-2},$
where the first inequality follows from Jensen’s inequality. The case $N=1$ is clear. It follows by Lemma 45 in Appendix D, that for all $\lambda\geq 1$
@@
By the lower bound in Lemma 49 we have for all $x\in\mathbb{S}$ and $\lambda\geq s$ a constant $c=\frac{s-1}{2\hat{w}+s-1}\in(0,1)$ such that
@@ Appendix F Cost function
$u(\lambda)-c(\lambda)=\frac{d(a+b\lambda)}{\lambda-1}\xrightarrow{\lambda\to\infty}bd,$
@@
Lemma 52.
Let Assumption (A1) hold, $f\in L_{0,+}^{2}(\pi)$ and $\lambda_{1},\lambda_{2}\in(1,2]$, $\lambda_{1}<\lambda_{2}$, then
$\lambda_{1}\cdot\mathrm{var}(P_{\lambda_{1}},f)-\lambda_{2}\cdot\mathrm{var}(P_{\lambda_{2}},f)>0.$
Proof.
Let $g\in L_{0}^{2}(\pi)$, then
$\mathcal{E}_{\lambda_{1}}(g)$ $=(\lambda_{1}-\lambda_{2})\mathcal{E}_{2}(g)+\mathcal{E}_{\lambda_{2}}(g).$
By Lemma 26, we can set $h_{\lambda_{2}}=(I-P_{\lambda_{2}})^{-1}f$, for which
$\mathrm{var}(P_{\lambda_{2}},f)=2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\langle f|f\rangle_{\pi},\qquad\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})=\langle h_{\lambda_{2}}|f\rangle_{\pi}=\langle f|h_{\lambda_{2}}\rangle_{\pi},$
and by Lemma 60 in Appendix H
$\mathrm{var}(P_{\lambda_{1}},f)=2\sup_{g\in L_{0}^{2}(\pi)}[2\langle f|g\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(g)]-\langle f|f\rangle_{\pi}\geq 2[2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(h_{\lambda_{2}})]-\langle f|f\rangle_{\pi}.$
Therefore
$\lambda_{1}\cdot\mathrm{var}(P_{\lambda_{1}},f)-\lambda_{2}\cdot\mathrm{var}(P_{\lambda_{2}},f)$
$\quad\geq\lambda_{1}2[2\langle f|h_{\lambda_{2}}\rangle_{\pi}-\mathcal{E}_{\lambda_{1}}(h_{\lambda_{2}})]-\lambda_{2}2\langle f|h_{\lambda_{2}}\rangle_{\pi}+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad=\lambda_{1}2[2\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})-(\lambda_{1}-\lambda_{2})\mathcal{E}_{2}(h_{\lambda_{2}})-\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})]-2\lambda_{2}\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad=\lambda_{1}2(\lambda_{2}-\lambda_{1})\mathcal{E}_{2}(h_{\lambda_{2}})-2(\lambda_{2}-\lambda_{1})\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad\geq 2(\lambda_{2}-\lambda_{1})[\mathcal{E}_{2}(h_{\lambda_{2}})-\mathcal{E}_{\lambda_{2}}(h_{\lambda_{2}})]+(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}$
$\quad\geq(\lambda_{2}-\lambda_{1})\langle f|f\rangle_{\pi}>0,$
where the second last inequality follows from Theorem 20 and the fact that $\lambda_{1}>1$. ∎
For Algorithm 2, derivatives of holding probability and $\psi$ (given in (14)) are defined for non-integer $\lambda>1$,
$\frac{\partial}{\partial\lambda}P_{\lambda}(x,\{x\})$ $=P_{\lfloor\lambda+1\rfloor}(x,\{x\})-P_{\lfloor\lambda\rfloor}(x,\{x\}),\qquad\psi^{\prime}(\lambda)=\psi(\lfloor\lambda+1\rfloor)-\psi(\lfloor\lambda\rfloor).$
Lemma 53.
Let $g,h:(1,\infty)\to\mathbb{R}$,
$g(\lambda)=\lambda\frac{1+\epsilon(\lambda)}{1-\epsilon(\lambda)},\qquad h(\lambda)=\lambda\frac{1+\psi(\lambda)}{1-\psi(\lambda)},$
then $g$ is strictly decreasing and strictly convex when $\lambda\in(1,2]$, same holds for $h$ if $(\mathbb{S})>1$.
Proof.
We have for $\lambda>1$, that $0<\epsilon(\lambda)<1$, and $\epsilon$ is decreasing by Theorem 21. Same properties hold for $\psi$ when $(\mathbb{S})>1$ by Lemma 25. We have for $\lambda\in(1,2)$, that
$g^{\prime}(\lambda)=\frac{2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}}{(1-\epsilon(\lambda))^{2}},$
where
$2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}$ $=2\lambda\big(\epsilon(2)-1\big)+1-\big((2-\lambda)+(\lambda-1)\epsilon(2)\big)^{2}$
$=2\lambda\big(\epsilon(2)-1\big)+2(\lambda-1)\big(1-\epsilon(2)\big)-(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}$
$=-2\big(1-\epsilon(2)\big)-(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}<0,$
and
$g^{\prime\prime}(\lambda)$ $=\frac{(1-\epsilon(\lambda))[2\epsilon^{\prime}(\lambda)-2\epsilon(\lambda)\epsilon^{\prime}(\lambda)]+2\epsilon^{\prime}(\lambda)[2\lambda\epsilon^{\prime}(\lambda)+1-\epsilon(\lambda)^{2}]}{(1-\epsilon(\lambda))^{3}}$
$=-\frac{2\epsilon^{\prime}(\lambda)\big[-(1-\epsilon(\lambda))^{2}+2\big(1-\epsilon(2)\big)+(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}\big]}{(1-\epsilon(\lambda))^{3}}$
$\geq-\frac{2\epsilon^{\prime}(\lambda)\big[\big(1-\epsilon(2)\big)+(\lambda-1)^{2}\big(1-\epsilon(2)\big)^{2}\big]}{(1-\epsilon(\lambda))^{3}}>0.$
Since $g$ is continuous, we get the claim. The proof is similar for $h$. ∎
When we combine the affine cost function with functions in (15), we obtain the following proposition:
Proposition 54.
Let $c:(1,\infty)\to\mathbb{R}$, $c(\lambda)=a+b\lambda$, where $a\geq 0$ and $b>0$. Let Assumption (A1) hold and $f\in L_{0,+}^{2}(\pi)$, then $\lambda\mapsto c(\lambda)V_{f}(\lambda)$, $\lambda\mapsto c(\lambda)G_{f}(\lambda)$ and $\lambda\mapsto c(\lambda)H_{f}(\lambda)$ are strictly decreasing when $\lambda\in(1,2]$.
Proof.
Follows from Theorem 27, Lemma 24 and 25 combined with Lemma 52 and 53. ∎
@@ Appendix G Finite discrete case for i-SIR
Estimating the asymptotic variance of i-SIR can be challenging. In the finite discrete case asymptotic variance can be solved via the formula in Lemma 39 in Appendix B, since the transition probability matrix $P_{\lambda}$ as well as the eigenvalues and eigenvectors can be solved, even tough this can be extremely slow for large $\lambda$ when $\mathbb{S}$ has many states. However, in the finite case transition probability matrix $P_{N}$ can be expressed with the help of multinomial distribution:
@@
Since $(X_{i})$ is Markov chain we obtain the case for $U=\emptyset$. The case $U\neq\emptyset$ follows similarly when we have instead $Q=(q(\{s_{1}\}),\dots,q(\{s_{n}\}),q(U))$, $v\in\mathbb{R}^{n+1}$,
@@
If $(\mathbb{S})=2$, i.e. $\mathbb{S}=\{s_{1},s_{2}\}$, then it holds for $f\in L_{0,+}^{2}(\pi)$ that
@@
We have $P_{\lambda}$ is $\pi$-reversible by Lemma 26. We can solve eigenvalues $\lambda_{1}=1$, $\lambda_{2}=P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})-1$ and eigenvectors $v_{1}=(1,1)$, $v_{2}=(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}},\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}})$ to match conditions in Lemma 39 in Appendix B and thus
$\mathrm{var}(P_{\lambda},f)$ $=\langle v_{2}|f\rangle_{\pi}^{2}\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-(P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\}))}$
@@
$\langle v_{2}|f\rangle_{\pi}^{2}$ $=\bigg(-\sqrt{\frac{\pi(\{s_{2}\})}{\pi(\{s_{1}\})}}f(s_{1})\pi(\{s_{1}\})+\sqrt{\frac{\pi(\{s_{1}\})}{\pi(\{s_{2}\})}}f(s_{2})\pi(\{s_{2}\})\bigg)^{2}$
@@
$-1+\sum_{i=1}^{2}\pi(\{s_{i}\})P_{\lambda}(s_{i},\{s_{i}\})=-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})$
$\quad=-[\pi(\{s_{1}\})+\pi(\{s_{2}\})][\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{2}\})+\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{1}\})]$
@@
$H_{f}(\lambda)$ $=\frac{1-2\pi(\{s_{1}\})^{2}-2\pi(\{s_{2}\})^{2}+\sum_{i=1}^{2}\pi(\{s_{i}\})P_{\lambda}(s_{i},\{s_{i}\})]}{1-\pi(\{s_{1}\})P_{\lambda}(s_{1},\{s_{1}\})-\pi(\{s_{2}\})P_{\lambda}(s_{2},\{s_{2}\})}\mathrm{var}_{\pi}(f)$
$=\frac{2-P_{\lambda}(s_{1},\{s_{2}\})-P_{\lambda}(s_{2},\{s_{1}\})]}{P_{\lambda}(s_{1},\{s_{2}\})+P_{\lambda}(s_{2},\{s_{1}\})}\mathrm{var}_{\pi}(f)=\frac{P_{\lambda}(s_{1},\{s_{1}\})+P_{\lambda}(s_{2},\{s_{2}\})}{2-P_{\lambda}(s_{1},\{s_{1}\})-P_{\lambda}(s_{2},\{s_{2}\})]}\mathrm{var}_{\pi}(f)$